# Learning to detect continuous gravitational waves: an open data-analysis competition

### v1

@@ 4 Competition results
The challenge actually makes use of two leaderboards, namely public leaderboard, which is visible throughout the duration of the competition, using a small fraction of the competition’s dataset to compute the AUC score, and the private leaderboard, which computes the AUC score using the complete dataset. This design allows participants to obtain a continuous assessment of their performance against other competitors, which is beneficial for a rapid iteration of the solution’s design, and minimises the chances of winning solutions ovefitting to a specific dataset. Additionally, the use of a random subset of the dataset minimises the probability of competitors successfully inferring the true classification of test data samples through repeated sumbmission [98, 99, 100].
@@ 4.1.1 Track-Statistic
The first class contains solutions implementing a track-based detection statistic similar to those used in blind CW searches with short coherence times [26, 27, 101, 102, 28, 29]. Specifically, these methods assume a closed-form frequency evolution (i.e. a “frequency track” in the spectrogram) parameterised by a small number of parameters (such as frequency, spindown, and sky position) along which a detection statistic is accumulated. The search then consists in finding the parameters that maximise said detection statistic. This class includes the teams JunKoda [103] (1st), PreferredWave [104] (2nd), HiddenNeuralLayers [105] (5th), and Shun_PI [106] (6th). These methods present two main degrees of freedom, namely the detection statistic and strategy to cover the parameter space, which we summarize in Table 3. To address non-Gaussianities arising in real-data samples, all teams masked frequency bands in the spectrogram containing large deviations from Gaussian noise.
@@
JunKoda (1st) Yes Maximum Dirichlet-weighed power [107] Square grid ($3.5\times 10^{8}$ templates)
@@
HiddenNeuralLayers (5th) No Maximum power Differential evolution (100 steps) [108]
Shun_PI (6th) No Maximum power Simulated-annealing [109].
@@
The relative placement of these four teams in the leaderboard is consistent with their strategies. HiddenNeuralLayers and Shun_PL place templates in a data driven manner. While such a strategy has been proven effective in local analyses [7, 32, 33, 34], the breadth of the CW parameter-space makes them unsuitable for broad searches such as the ones here considered [7], hence their relatively lower placement. The difference between JunKoda and PreferredWave, on the other hand, can be ascribed to the significant difference in the number of evaluated templates ($10^{8}$ in JunKoda versus $10^{6}$ in PreferredWave) and the fact that JunKoda takes the maximum rather than the average power in the template bank [29].
@@ 4.1.2 Track-Finding
The second class introduces an extra degree of freedom with respect to Sec. 4.1.1, namely that the frequency evolution of the signal is not parameterised but inferred from the data. This approach is used by team SpaceCoders (4th), and bears some similarities to CW searches based on the Viterbi algorithm [53].
Specifically, the method uses dynamic programming to find a track maximising the accumulated SFT power. The track is constrained to contain a single inflection point, given the duration of the dataset and the expected signal morphology. Upon identifying a track, the signal’s parameters are estimated using a parametric fit in terms of frequency, spindown, and sky position. The optimisation is carried out using the Nelder-Mead algorithm [110], in a similar manner to the methods presented in Sec. 4.1.1.
The relative loss of SpaceCoders with respect to other track-based methods can be justified as follows: First, the number of tracks explored by the SpaceCoders search is exponentially greater than a closed-form track search [111, 53, 5]; this increases the expected background distribution [112] and, consequently, lowers the detection probability at a fixed amplitude. Second, the track partially relies on the signal displaying strong features within short timescales to clearly identify it. Due to the idiosyncrasy of CWs, this is only true for relatively strong signals.
@@ 4.1.3 Machine-Learning
The final class includes solutions using machine-learning strategies; that is, search pipelines in which some component is trained before producing a submission. This class comprises teams BearWaves [113] (3rd), KDL [^1] No available information about this team (7th), MCDigital [114] (8th), yu4u [115] (9th), and anonamename [116] (10th).
anonamename implemented a classifier based on a Convolutional Neural Network (CNN). Data samples are pre-processed following [116] and two outputs are produced: a detection statistic, and the estimated location of the signal in the spectrogram as a mask. To account for real data, line-like features were generated and included during training.
yu4u [115] implemented a similar strategy using a U-net architecture to estimate the location of the signal in the spectrogram and compute a detection statistic. As in the previous case, real data noise features were addressed by generating data containing line-like features during training.
MC Digital [114] combines two detection statistics using a CNN and the $\mathcal{F}$-statistic as computed by PyFstat [81]. Gaussian and real-data are addressed separately by training two different networks. The combination of CNN and $\mathcal{F}$-statistic results is done manually: samples with high $\mathcal{F}$-statistic values are classified according to the $\mathcal{F}$-statistic; samples with low $\mathcal{F}$-statistic values are classified according to the CNN.
Finally, BearWaves [113] presented a more involved approach. First, training data was simulated on-the-fly using PyFstat to avoid overfitting. To simplify the task of combining data from two detectors, data samples were time-averaged to obtain a common set of timsetamps. Data samples were analyzed by an ensemble of neural networks applying different processing strategies.
Based on their relative leaderboard positions, image-based machine-learning strategies tend to deliver less sensitive pipelines compared to track-based statistics unless a significant amount of resources is invested in data pre- and post-processing, as shown by BearWaves [113].
@@ 4.2 Reference submission: SOAP
We now describe the application of SOAP [53, 54, 55] to the competition’s dataset. SOAP is a fast, model-agnostic blind CW search capable of detecting a broad class of CW sources, even if they are affected by complicated physics such as spin-wandering, glitches, or timing noise [9]. This robustness, however, limits the sensitivity of SOAP to well-modelled isolated NSs, as they correspond to a minuscule portion of the probed parameter space [5].
Contrary to most blind CW searches, which identify interesting parameter-space regions where a signal may be present [32], SOAP’s main search stage assigns detection statistics to frequency bands by returning the most-significant track as constructed by the Viterbi algorithm [117]. This is consistent with the submission format expected in the Kaggle competition.
The agnosticity of SOAP makes it an ideal pipeline to conduct a first comparison between Kaggle solutions and real CW searches. More concretely, we require Kaggle solutions to return a greater AUC score than that achieved by SOAP, as otherwise their design choices would make them less performant than a search with minimal assumptions.
The SOAP analysis begins 1800s SFT power spectra, which are summed over one-day periods (48 SFTs). This daily summation process effectively average out antenna pattern modulation and increasing the SNR in each frequency bin. The search uses the line-aware statistic [53], a multi-detector approach that computes the Bayesian odds ratio between the signal hypothesis and the noise hypothesis including spectral lines [28]. This statistic specifically penalises instrumental line-like features in spectrograms, helping to distinguish genuine astrophysical signals from detector artifacts.
The line aware statistic incorporates three tunable parameters: the width of the prior distribution on signal SNR $\rho_{\rm{signal}}$, the width of the prior distribution on line SNR $\rho_{\rm{line}}$, and the ratio of prior probabilities between noise and line models $p(\mathrm{line})/p(\mathrm{noise})$. Following the search in LIGO O3 data [13], these parameters are set to $\rho_{\rm{signal}}=4.0$, $\rho_{\rm{line}}=10.0$, and $p(\mathrm{line})/p(\mathrm{noise})=0.4$.
Additional tunable parameters include the transition matrix probabilities that guide the algorithm’s frequency tracking behavior. Using the two-detector configuration, the transition matrix is configured to constrain the algorithm to search identical frequency bins across both detectors. This results in a three-element transition matrix defining probabilities for frequency transitions $(+T_{\textrm{SFT}}^{-1},0,-T_{\textrm{SFT}}^{-1})$ with equal values matching those used in the O3 search [13].
For each frequency band in the test dataset described in Section 3, a value of the line-aware statistic is produced for each band.
@@ 4.3 Comparison
All solutions in the top 10 leaderboard produce an AUC above the value reported by the SOAP reference submission in Sec. 4.2, implying that all pipelines introduce cogent information into their assumptions when compared to an agnostic search.
The AUC score, however, aggregates results computed from the complete dataset across multiple false-positive probabilities. As a result, there is no single solution that dominates across all false-positive probabilities. Also, the fact that signals with different SNRs are combined together makes a direct comparison with standard CW search methods difficult.
@@ 5 Comparison to CW searches
Strictly speaking, $c_{t}$ and $n_{t}$ depend on the specific implementation of a CW search [6, 118, 24, 28, 25, 29]. We here instead take an educated guess based on the latest result in the literature. The latest GCT searches conducted in LIGO O2 and O3 data [86, 12] use $n_{\mathrm{t}}\sim 10^{12}-10^{15}$, depending on the frequency band. Similarly, the computing cost of the $\mathcal{F}$-statistic is about $c_{\mathrm{{t}}}\sim(10-100)\,\mathrm{ns}$ [22, 86, 12, 87]. Thus, $c_{\mathrm{t}}n_{\mathrm{t}}\sim(10^{4}-10^{7})\,\mathrm{s}$.
@@
Figure 5: False dismissal probability as a function of false-positive probability for a restricted range of SNR. The upper row shows the results using Gaussian noise samples, while the lower row uses only real-data samples. The shaded regions show binominal errorbars.
@@
For the GCT range ($\rho\in[50,60]$), in Gaussian noise, JunKoda, Preferred Wave and Space Coders reach $p_{\mathrm{FP}}\sim 10^{-3}$ at the specified false dismissal probability level. The next relevant team is yu4yu (9th), at $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. The remaining solutions about $p_{\mathrm{FP}}\sim 10^{-1}$.
Despite the increased uncertainty, the situation in real-data samples changes slightly: JunKoda stays at $p_{\mathrm{FP}}\sim 10^{-3}$; Preferred Wave worsens up to $p_{\mathrm{FP}}\sim 7\times 10^{-3}$, and Space Coders loses an order of magnitude up to $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. yu4u maintains a comparable performance to the Gaussian case. The relative loss of PreferredWave with respect to JunKoda can be ascribed to the use of average power in a template bank: In non-Gaussian data, the average template-bank power tends to rise due to the presence of extreme values in the background, even if most of the problematic bins have been removed. Similarly, the loss of Space Coders can be related to the fact that their method builds-up a frequency-evolution track in a data-driven manner, an thus can get biased due to local maxima in the spectrogram. Finally, we highlight the performance of Hidden Neural Layers (5th), which reaches $p_{\mathrm{FP}}\sim 10^{-3}$ surpasing all other teams except JunKoda. This is likely a consequence of their approach to isolate and remove potential glitches [105].
@@ 6 Discussion
CW signals are one of the potential next discoveries of gravitational-wave astronomy. Currently their detection is complicated by our ignorance on the specific emission mechanism at play, which gives rise to a broad parameter-space that requires significant computing budgets to be effectively covered. Fundamentally, the sensitivity of CW searches for unknown sources is limited by the available computing [6, 24, 5].
In this work, we explored the use of data-analysis competitions to discover new search strategies. Specifically, we framed CW searches as a classification problem and deployed a Kaggle competition to find the most effective strategy to identify data samples harbouring simulated CW signals. Kaggle [73] is a online data-analysis competition platform powered by Google hosting a broad community of data-analysis practitioners with a wide variety of backgrounds. By tapping into a new, broad community of experts, we intended to asses the performance of data-analysis methods from other field when applied to the problem of detecting CW signals. This approach is similar to mock data challenges previously developed for other gravitational-wave use cases [57, 58, 60, 59].
@@
Overall, the competition was dominated by pipelines based on semicoherent matched filtering, in a similar fashion to [26, 27]. The proposed solutions, however, introduced several improvements with respect to well established methods.
First, the evaluation of detection statistics was vectorised over multiple waveforms using GPUs without using any signal approximation. This contrasts with most GW GPU-accelerated methods so far, which either parallelize the evaluation of a single template [22, 119] or compute an approximated version of the detection statistic [28, 120]. The potential improvement in computational cost and complexity of search pipelines inspired the development of fasttracks [29], a unified, GPU-accelerated engine to compute CW detection statistics using SFTs for general CW waveform models.
Second, the winning solution [103] made use of alternative filtering strategies, similar to those previously outlined in Ref. [107]. While these specific statistics have not been applied to short-coherence CW searches, their efficiency allowed us to extend the use-case of efficient SFT-based detection statistics to the search for binary neutron stars and massive black-hole binaries [121] in next-generation detectors (Einstein Telescope [122], Cosmic Explorer [123], LISA [124]).
Solutions accounting for the presence of non-Gaussian data did so in a relatively simple manner, either by masking problematic regions in the spectrogram [103, 104, 105, 106] or by generating non-Gaussian training data [113, 114, 115, 116]. This is comparable to the practical approach taken in CW searches, where non-Gaussianities can be identified as sudden spikes in the accumulated power [28, 29].
@@
The relevance of approaches based on alternative techniques, such as machine learning, becomes apparent when analyzing specific subsets of SNR. For example, for moderately strong signals $\rho\in[50,60]$, we find that U-net architectures [115] are able to compete with track-based statistics and show a negligible performance degradation when applied to non-Gaussian data.
Despite the limited duration of this challenge, the obtained solutions have the potential to affect future CW searches not only by improving the efficiency and deployment of track-based searches [29, 121], but also by spurring approaches prescinding from template banks all together. This latter point is particularly interesting, as the computing cost of a CW search is dominated by the evaluation of statistics on a template bank to numerically marginalise with respect to the signal parameters to compute a Bayes factor [76, 77, 78]; this Kaggle competition, on the other hand, allows for pipelines operating directly on data samples, so that the method itself constitutes an approximation of the marginalised Bayes factor. The solutions here collected, together with those proposed from within the community [48, 125, 54, 49, 55, 52, 50], bear witness of the potential of alternative approaches to revolutionize the search for blind CW sources.
Although the dataset is shorter than typical production searches, the chosen duration (4 months) remains a valid regime for candidate generation. Given the expected cadence of forthcoming LVK observing runs, these results support a strategy based on analysing each run shallowly to produce candidate lists, with significance subsequently accumulated across runs—akin to the “fresh-data mode” discussed in Ref. [126]. Such implementation will be presented elsewhere.
This Kaggle competition is the first public, large-scale exposure of the CW detection problem to a wide community of data-analysis practitioners, and a logical step beyond established closed-door mock-data challenges [57, 58, 60, 59] to foster cross-pollination across fields. Its impact is already evident in new GPU-accelerated methods [29] and in applying CW-inspired detection statistics to other domains [121]. The materials developed for this competition, including tutorials for generating and pre-processing CW data with PyFstat [81], and the dataset itself [61], are publicly available to support further development of alternative methods aimed at the first CW detection.

### v2

@@ 4 Competition results
The challenge actually makes use of two leaderboards, namely public leaderboard, which is visible throughout the duration of the competition, using a small fraction of the competition’s dataset to compute the AUC score, and the private leaderboard, which computes the AUC score using the complete dataset. This design allows participants to obtain a continuous assessment of their performance against other competitors, which is beneficial for a rapid iteration of the solution’s design, and minimises the chances of winning solutions ovefitting to a specific dataset. Additionally, the use of a random subset of the dataset minimises the probability of competitors successfully inferring the true classification of test data samples through repeated sumbmission [Whitehill15, Whitehill17climb, Whitehill17].
@@ 4.1.1 Track-Statistic
The first class contains solutions implementing a track-based detection statistic similar to those used in blind CW searches with short coherence times. Specifically, these methods assume a closed-form frequency evolution (i.e. a “frequency track” in the spectrogram) parameterised by a small number of parameters (such as frequency, spindown, and sky position) along which a detection statistic is accumulated. The search then consists in finding the parameters that maximise said detection statistic. This class includes the teams JunKoda [first_team] (1st), PreferredWave [second_team] (2nd), HiddenNeuralLayers [fifth_team] (5th), and Shun_PI [sixth_team] (6th). These methods present two main degrees of freedom, namely the detection statistic and strategy to cover the parameter space, which we summarize in Table 3. To address non-Gaussianities arising in real-data samples, all teams masked frequency bands in the spectrogram containing large deviations from Gaussian noise.
@@
JunKoda (1st) Yes Maximum Dirichlet-weighed power [Allen:2002bp] Square grid ($3.5\times 10^{8}$ templates)
@@
HiddenNeuralLayers (5th) No Maximum power Differential evolution (100 steps) [DiffEv]
Shun_PI (6th) No Maximum power Simulated-annealing [SimAn].
@@
The relative placement of these four teams in the leaderboard is consistent with their strategies. HiddenNeuralLayers and Shun_PI place templates in a data driven manner. While such a strategy has been proven effective in local analyses [Ashton:2017wui, Tenorio:2021njf, Covas:2024pam, Mirasola:2024lcq], the breadth of the CW parameter-space makes them unsuitable for broad searches such as the ones here considered [Ashton:2017wui], hence their relatively lower placement. The difference between JunKoda and PreferredWave, on the other hand, can be ascribed to the significant difference in the number of evaluated templates ($10^{8}$ in JunKoda versus $10^{6}$ in PreferredWave) and the fact that JunKoda takes the maximum rather than the average power in the template bank [Tenorio:2024jgc].
@@ 4.1.2 Track-Finding
The second class introduces an extra degree of freedom with respect to Sec. 4.1.1, namely that the frequency evolution of the signal is not parameterised but inferred from the data. This approach is used by team SpaceCoders (4th), and bears some similarities to CW searches based on the Viterbi algorithm [soap].
Specifically, the method uses dynamic programming to find a track maximising the accumulated SFT power. The track is constrained to contain a single inflection point, given the duration of the dataset and the expected signal morphology. Upon identifying a track, the signal’s parameters are estimated using a parametric fit in terms of frequency, spindown, and sky position. The optimisation is carried out using the Nelder-Mead algorithm [nelder_mead], in a similar manner to the methods presented in Sec. 4.1.1.
The relative loss of SpaceCoders with respect to other track-based methods can be justified as follows: First, the number of tracks explored by the SpaceCoders search is exponentially greater than a closed-form track search [Wette:2023dom]; this shifts the expected background distribution upwards [Tenorio:2021wad] and, consequently, lowers the detection probability at a fixed amplitude. Second, the track partially relies on the signal displaying strong features within short timescales to clearly identify it. Due to the idiosyncrasy of CWs, this is only true for relatively strong signals.
@@ 4.1.3 Machine-Learning
The final class includes solutions using machine-learning strategies; that is, search pipelines in which some component is trained before producing a submission. This class comprises teams BearWaves [third_team] (3rd), KDL [^1] No available information about this team (7th), MCDigital [eighth_team] (8th), yu4u [nineth_team] (9th), and anonamename [tenth_team] (10th).
anonamename implemented a classifier based on a Convolutional Neural Network (CNN). Data samples are pre-processed following [tenth_team] and two outputs are produced: a detection statistic, and the estimated location of the signal in the spectrogram as a mask. To account for real data, line-like features were generated and included during training.
yu4u [nineth_team] implemented a similar strategy using a U-net architecture to estimate the location of the signal in the spectrogram and compute a detection statistic. As in the previous case, real data noise features were addressed by generating data containing line-like features during training.
MC Digital [eighth_team] combines two detection statistics using a CNN and the $\mathcal{F}$-statistic as computed by PyFstat [pyfstat]. Gaussian and real-data are addressed separately by training two different networks. The combination of CNN and $\mathcal{F}$-statistic results is done manually: samples with high $\mathcal{F}$-statistic values are classified according to the $\mathcal{F}$-statistic; samples with low $\mathcal{F}$-statistic values are classified according to the CNN.
Finally, BearWaves [third_team] presented a more involved approach. First, training data was simulated on-the-fly using PyFstat to avoid overfitting. To simplify the task of combining data from two detectors, data samples were time-averaged to obtain a common set of timsetamps. Data samples were analyzed by an ensemble of neural networks applying different processing strategies.
Based on their relative leaderboard positions, image-based machine-learning strategies tend to deliver less sensitive pipelines compared to track-based statistics unless a significant amount of resources is invested in data pre- and post-processing, as shown by BearWaves [third_team].
@@ 4.2 Reference submission: SOAP
We now describe the application of SOAP [soap] to the competition’s dataset. SOAP is a fast, model-agnostic blind CW search capable of detecting a broad class of CW sources, even if they are affected by complicated physics such as spin-wandering, glitches, or timing noise [Carlin:2025sxm]. This robustness, however, limits the sensitivity of SOAP to well-modelled isolated NSs due the breadth of its parameter space [Wette:2023dom].
Contrary to most blind CW searches, which identify interesting parameter‑space templates or regions where a signal may be present [Tenorio:2021njf], SOAP’s main search stage assigns detection statistics to frequency bands by returning the most significant track constructed via the Viterbi algorithm [viterbi]. This matches the submission format of the Kaggle competition and makes SOAP a natural baseline for comparing against Kaggle solutions.
The SOAP analysis begins 1800s SFT power spectra, which are summed over one-day periods (48 SFTs). This daily summation process effectively average out antenna pattern modulation and increasing the SNR in each frequency bin. The search uses the line-aware statistic [soap], a multi-detector approach that computes the Bayesian odds ratio between the signal hypothesis and the noise hypothesis including spectral lines [Covas:2019jqa]. This statistic specifically penalises instrumental line-like features in spectrograms, helping to distinguish genuine astrophysical signals from detector artifacts. We use the same SOAP configuration as that used in [KAGRA:2022dwb]. For each frequency band in the test dataset described in Sec. 3, a value of the line-aware statistic is produced for each band.
Note that SOAP’s signal model only requires continuity in the frequency-evolution track across consecutive days. No restrictions are set on the coherence of the signal, or the specific functional form of the frequency and amplitude modulation due to the detector’s motion. This freedom gives rise to the large parameter space covered by this search. As a result, if a Kaggle solution surpasses the performance of SOAP, it suggests an enhanced capability to identify structures in the data that are characteristic of CW signals, likely because the method incorporates more informative signal priors in its design.
@@ 4.3 Comparison
All solutions in the top 10 leaderboard produce an AUC above the value reported by the SOAP reference submission in Sec. 4.2. This indicates that all pipelines introduce cogent information into their assumptions when compared to an agnostic search, as previously discussed.
The AUC score, however, aggregates results computed from the complete dataset, which contains signals with different SNRS, across multiple false-positive probabilities. As a result, there is no single solution that dominates across all false-positive probabilities. To more directly compare to standard CW search methods, we present results as a function of SNR in Sec. 5.
@@ 5 Comparison to CW searches
Strictly speaking, $c_{t}$ and $n_{t}$ depend on the specific implementation of a CW search [Wette:2023dom]. We here instead take an educated guess based on the latest result in the literature. The latest GCT searches conducted in LIGO O2 and O3 data [eah_O2, Steltner:2023cfk] use $n_{\mathrm{t}}\sim 10^{12}-10^{15}$, depending on the frequency band. Similarly, the computing cost of the $\mathcal{F}$-statistic is about $c_{\mathrm{{t}}}\sim(10-100)\,\mathrm{ns}$ [Dunn:2022gai, eah_O2, Steltner:2023cfk, Wette:2021tbv]. Thus, $c_{\mathrm{t}}n_{\mathrm{t}}\sim(10^{4}-10^{7})\,\mathrm{s}$.
@@
Figure 6: False dismissal probability as a function of false-positive probability for a restricted range of SNR. The upper row shows the results using Gaussian noise samples, while the lower row uses only real-data samples. The shaded regions show binominal errorbars $\delta p_{\mathrm{FD}}=\sqrt{p_{\mathrm{FD}}(1-p_{\mathrm{FD}})/N_{\mathrm{s}}}$, where $N_{\mathrm{s}}$ is the number of signals in said SNR range.
@@
For the GCT range ($\rho\in[50,60]$), in Gaussian noise, JunKoda, Preferred Wave and Space Coders reach $p_{\mathrm{FP}}\sim 10^{-3}$ at the specified false dismissal probability level. The next relevant team is yu4u (9th), at $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. The remaining solutions about $p_{\mathrm{FP}}\sim 10^{-1}$.
Despite the increased uncertainty, the situation in real-data samples changes slightly: JunKoda stays at $p_{\mathrm{FP}}\sim 10^{-3}$; Preferred Wave worsens up to $p_{\mathrm{FP}}\sim 7\times 10^{-3}$, and Space Coders loses an order of magnitude up to $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. yu4u maintains a comparable performance to the Gaussian case. The relative loss of PreferredWave with respect to JunKoda can be ascribed to the use of average power in a template bank: In non-Gaussian data, the average template-bank power tends to rise due to the presence of extreme values in the background, even if most of the problematic bins have been removed. Similarly, the loss of Space Coders can be related to the fact that their method builds-up a frequency-evolution track in a data-driven manner, an thus can get biased due to local maxima in the spectrogram. Finally, we highlight the performance of Hidden Neural Layers (5th), which reaches $p_{\mathrm{FP}}\sim 10^{-3}$ surpassing all other teams except JunKoda. This is likely a consequence of their approach to isolate and remove potential glitches [fifth_team].
@@ 6 Discussion
CW signals are one of the potential next discoveries of gravitational-wave astronomy. Currently their detection is complicated by our ignorance on the specific emission mechanism at play, which gives rise to a broad parameter-space that requires significant computing budgets to be effectively covered. Fundamentally, the sensitivity of CW searches for unknown sources is limited by the available computing [Prix:2012yu, Wette:2018bhc, Wette:2023dom].
In this work, we explored the use of data-analysis competitions to discover new search strategies. Specifically, we framed CW searches as a classification problem and deployed a Kaggle competition to find the most effective strategy to identify data samples harbouring simulated CW signals. Kaggle [kaggle] is a online data-analysis competition platform powered by Google hosting a broad community of data-analysis practitioners with a wide variety of backgrounds. By tapping into a new, broad community of experts, we intended to asses the performance of data-analysis methods from other field when applied to the problem of detecting CW signals. This approach is similar to mock data challenges previously developed for other gravitational-wave use cases [Messenger:2015tga, Walsh:2016hyc, Baghi:2022ucj, Schafer:2022dxv].
@@
Overall, the competition was dominated by pipelines based on semicoherent matched filtering, in a similar fashion to [Krishnan:2004sv, 2014PhRvD..90d2002A]. The proposed solutions, however, introduced several improvements with respect to well established methods.
First, the evaluation of detection statistics was vectorised over multiple waveforms using GPUs without using any signal approximation. This contrasts with most GW GPU-accelerated methods so far, which either parallelize the evaluation of a single template [Dunn:2022gai, Garcia-Quiros:2025usi] or compute an approximated version of the detection statistic [Covas:2019jqa, Rosa:2021ptb]. The potential improvement in computational cost and complexity of search pipelines inspired the development of fasttracks [Tenorio:2024jgc], a unified, GPU-accelerated engine to compute CW detection statistics using SFTs for general CW waveform models.
Second, the winning solution [first_team] made use of alternative filtering strategies, similar to those previously outlined in Ref. [Allen:2002bp]. While these specific statistics have not been applied to short-coherence CW searches, their efficiency allowed us to extend the use-case of efficient SFT-based detection statistics to the search for binary neutron stars and massive black-hole binaries [Tenorio:2025gci] in next-generation ground-based detectors (Einstein Telescope [Abac:2025saz], Cosmic Explorer [Reitze:2019iox]). Similar strategies have been proposed to analyze GW signals in the LISA [LISA:2024hlh] frequency band, such as galactic binaries [Prix:2007zh, Blaut:2009si], which are modelled in a similar way to LIGO CW sources, and more recently a broader class of systems such as stellar-mass binary black holes [Bandopadhyay:2025fyx] and extreme mass-ratio inspirals [Speri:2025ucn].
Solutions accounting for the presence of non-Gaussian data did so in a relatively simple manner, either by masking problematic regions in the spectrogram [first_team, second_team, fifth_team, sixth_team] or by generating non-Gaussian training data [third_team, eighth_team, nineth_team, tenth_team]. This is comparable to the practical approach taken in CW searches, where non-Gaussianities can be identified as sudden spikes in the accumulated power [Covas:2019jqa, Tenorio:2024jgc].
@@
The relevance of approaches based on alternative techniques, such as machine learning, becomes apparent when analyzing specific subsets of SNR. For example, for moderately strong signals $\rho\in[50,60]$, we find that U-net architectures [nineth_team] are able to compete with track-based statistics and show a negligible performance degradation when applied to non-Gaussian data.
Despite the limited duration of this challenge, the obtained solutions have the potential to affect future CW searches not only by improving the efficiency and deployment of track-based searches [Tenorio:2024jgc, Tenorio:2025gci], but also by spurring approaches prescinding from template banks all together. This latter point is particularly interesting, as the computing cost of a CW search is dominated by the evaluation of statistics on a template bank to numerically marginalise with respect to the signal parameters to compute a Bayes factor [Prix:2009tq, Searle:2008jv, Prix:2011qv]; this Kaggle competition, on the other hand, allows for pipelines operating directly on data samples, so that the method itself constitutes an approximation of the marginalised Bayes factor. The solutions here collected, together with those proposed from within the community [Dreissigacker:2019edy, Miller:2019jtp, Bayley:2020zfa, Yamamoto:2020pus, Bayley:2022hkz, Modafferi:2023nzt, Joshi:2025xdz], bear witness of the potential of alternative approaches to revolutionize the search for blind CW sources.
Although the dataset is shorter than typical production searches, the chosen duration (4 months) remains a valid regime for candidate generation. Given the expected cadence of forthcoming LVK observing runs, these results support a strategy based on analysing each run shallowly to produce candidate lists, with significance subsequently accumulated across runs, akin to the “fresh-data mode” discussed in Ref. [Cutler:2005pn]. Such implementation will be presented elsewhere.
This Kaggle competition is the first public, large‑scale exposure of the CW detection problem to a wide community of data‑analysis practitioners, and a logical step beyond established closed‑door mock‑data challenges [Messenger:2015tga, Walsh:2016hyc, Baghi:2022ucj, Schafer:2022dxv] to foster cross‑pollination across fields. Its impact is already evident in new GPU‑accelerated methods [Tenorio:2024jgc] and in applying CW‑inspired detection statistics to other domains, such as compact binary coalescenses in next-generation GW detectors [Tenorio:2025gci, Speri:2025ucn, Bandopadhyay:2025fyx]. The materials developed for this competition, including tutorials for generating and pre‑processing CW data with PyFstat [pyfstat], and the dataset itself [dataset], are publicly available to support further development of alternative methods aimed at the first CW detection.